# Nested subspace learning with flags

### v1

@@ 4.3.1 Application of the flag trick to trace ratio problems
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}}B\right)}.$ (16)
@@
The flag trick applied to the TR problem (16) reads
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmax}}\,\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)}.$ (17)
and is equivalent to the following optimization problem:
$\underset{U_{1:d}\in\operatorname{St}(p,q)}{\operatorname{argmax}}\,\frac{\sum_{k=1}^{d}(d-(k-1))\operatorname{tr}\left({U_{k}}^{\top}A{U_{k}}\right)}{\sum_{l=1}^{d}(d-(l-1))\operatorname{tr}\left({U_{l}}^{\top}B{U_{l}}\right)}.$ (18)
Proof  The proof is given in Appendix (Appendix B).
Equation (18) tells us several things. First, the subspaces $\leavevmode\nobreak\ {\operatorname{Span}(U_{1})\perp\dots\perp\operatorname{Span}(U_{d})}$ are weighted decreasingly, which means that they have less and less importance with respect to the TR objective. Second, we can see that the nested trace ratio problem (17) somewhat maximizes the numerator $\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)$ while minimizing the denominator $\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)$. Both subproblems have an explicit solution corresponding to our nested PCA Theorem 4. Hence, one can naturally initialize the steepest descent algorithm with the $q$ highest eigenvalues of $A$ or the $q$ lowest eigenvalues of $B$ depending on the application. For instance, for LDA, initializing Algorithm 1 with the highest eigenvalues of $A$ would spread the classes far apart, while initializing it with the lowest eigenvalues of $B$ would concentrate the classes, which seems less desirable since we do not want the classes to concentrate at the same point.
@@ 4.3.2 Nestedness experiments for trace ratio problems
For all the experiments of this subsection, we consider the particular TR problem of LDA, although many other applications (marginal Fisher analysis (Yan et al., 2007), local discriminant embedding (Chen et al., 2005) etc.) could be investigated similarly.
First, we consider a synthetic dataset with five clusters. The ambient dimension is $p=3$ and the intrinsic dimensions that we try are $q_{1:2}=(1,2)$. We adopt a preprocessing strategy similar to Ngo et al. (2012): we first center the data, then run a PCA to reduce the dimension to $n-C$ (if $n-C<p$), then construct the LDA scatter matrices $A$ and $B$, then add a diagonal covariance regularization of $10^{-5}$ times their trace and finally normalize them to have unit trace. We run Algorithm 1 on Grassmann manifolds to solve the TR maximization problem (16), successively for $q_{1}=1$ and $q_{2}=2$. Then we plot the projections of the data points onto the optimal subspaces. We compare them to the nested projections onto the optimal flag output by running Algorithm 1 on $\operatorname{Fl}(3,(1,2))$ to solve (17). The results are shown in Figure 4.
@@
Figure 5: Illustration of the nestedness issue in linear discriminant analysis (trace ratio problem) on the digits dataset. For $q_{k}\in q_{1:d}:=(1,2,\dots,63)$, we solve the Grassmannian optimization problem (16) on $\operatorname{Gr}(64,q_{k})$ and plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ (left) and explained variances ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ (right) as a function of $k$. We compare those quantities to the ones obtained by solving the flag optimization problem (17). We can see that the Grassmann-based method is highly non-nested and even yields an extremely paradoxical non-increasing explained variance (cf. red circle on the right).
We can see that the Grassmann representations are non-nested while their flag counterparts perfectly capture the filtration of subspaces that best and best approximates the distribution while discriminating the classes. Even if the colors make us realize that the issue in this experiment for LDA is not much about the non-nestedness but rather about the rotation of the principal axes within the 2D subspace, we still have an important issue of consistency.
@@
Second, we consider the (full) handwritten digits dataset (Alpaydin and Kaynak, 1998). It contains $8\times 8$ pixels images of handwritten digits, from 00 to $9$, almost uniformly class-balanced. One has $n=1797$, $p=64$ and $C=10$. We run a steepest descent algorithm to solve the trace ratio problem (17). We choose the full signature $q_{1:63}=(1,2,\dots,63)$ and compare the output flag to the individual subspaces output by running optimization on $\operatorname{Gr}(p,q_{k})$ for $q_{k}\in q_{1:d}$. We plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ and the explained variance ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ as a function of the $k$. The results are illustrated in Figure 5. We see that the subspace angles are always positive and even very large sometimes with the LDA. Worst, the explained variance is not monotonous. This implies that we sometimes loose some information when increasing the dimension, which is extremely paradoxical.
@@
Third, we perform some classification experiments on the optimal subspaces for different datasets. For different datasets, we run the optimization problems on $\operatorname{Fl}(p,q_{1:d})$, then project the data onto the different subspaces in $\mathcal{S}_{1:d}^{*}$ and run a nearest neighbors classifier with $5$ neighbors. The predictions are then ensembled (cf. Algorithm 2) by weighted averaging, either with uniform weights or with weights minimizing the average cross-entropy:
$w_{1}^{*},\dots,w_{d}^{*}=\underset{\begin{subarray}{c}w_{k}\geq 0\\ \sum_{k=1}^{d}w_{k}=1\end{subarray}}{\operatorname{argmin}}\,-\frac{1}{nC}\sum_{i=1}^{n}\sum_{c=1}^{C}y_{ic}\ln\left(\sum_{k=1}^{d}w_{k}y_{kic}^{*}\right),$ (19)
@@
where $y_{kic}^{*}\in[0,1]$ is the predicted probability that $x_{i}\in\mathbb{R}^{p}$ belongs to class $c\in\{1\dots C\}$, by the classifier $g_{k}^{*}$ that is trained on $Z_{k}:={U_{k}^{*}}^{\top}X\in\mathbb{R}^{q_{k}\times n}$. One can show that the latter is a convex problem, which we optimize using the cvxpy Python package (Diamond and Boyd, 2016). We report the results in Table 1.
@@
Table 1: Classification results for the TR problem on real datasets. For each method (Gr: Grassmann optimization (16), Fl: flag optimization (16), Fl-U: flag optimization + uniform soft voting, Fl-W: flag optimization + optimal soft voting (19)), we give the cross-entropy of the projected-predictions with respect to the true labels.
@@
breast $569$ $30$ $(1,2,5)$ $0.0986$ $0.0978$ $0.0942$ $0.0915$ $(0.754,0,0.246)$
iris $150$ $4$ $(1,2,3)$ $0.0372$ $0.0441$ $0.0410$ $0.0368$ $(0.985,0,0.015)$
wine $178$ $13$ $(1,2,5)$ $0.0897$ $0.0800$ $0.1503$ $0.0677$ $(0,1,0)$
digits $1797$ $64$ $(1,2,5,10)$ $0.4507$ $0.4419$ $0.5645$ $0.4374$ $(0,0,0.239,0.761)$
The first example tells us that the optimal $5D$ subspace obtained by Grassmann optimization less discriminates the classes than the $5D$ subspace from the optimal flag. This may show that the flag takes into account some lower dimensional variability that enables to better discriminate the classes. We can also see that the uniform averaging of the predictors at different dimensions improves the classification. Finally, the optimal weights improve even more the classification and tell that the best discrimination happens by taking a soft blend of classifier at dimensions $1$ and $5$. Similar kinds of analyses can be made for the other examples.
@@ A Newton algorithm
In all the experiments of this paper, we use a steepest descent method on flag manifolds (Algorithm 1) to solve the flag problems. However, for the specific problem of TR (17), we believe that more adapted algorithms should be derived to take into account the specific form of the objective function, which is classically solved via a Newton-Lanczos method (Ngo et al., 2012). To that extent, we develop in the appendix (Appendix B) an extension of the baseline Newton-Lanczos algorithm for the flag-tricked problem (17). In short, the latter can be reformulated as a penalized optimization problem of the form $\operatorname{argmax}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho B)\right)}$, where $\rho$ is updated iteratively according to a Newton scheme. Once again, our central Theorem 4 enables to get explicit expressions for the iterations, which results without much difficulties in a Newton method, that is known to be much more efficient than first-order methods like the steepest descent.
@@
The classical trace ratio problems look for linear embeddings of the data. However, in most cases, the data follow a nonlinear distribution, which may cause linear dimension reduction methods to fail. The kernel trick (Hofmann et al., 2008) is a well-known method to embed nonlinear data into a linear space and fit linear machine learning methods. As a consequence, we propose in appendix (Appendix B) a kernelization of the trace ratio problem (17) in the same fashion as the one of the seminal graph embedding work (Yan et al., 2007). This is expected to yield much better embedding and classification results.
@@ 4.4 The flag trick for spectral clustering problems
Spectral clustering (Ng et al., 2001) is a baseline clustering technique. It can be applied to general cluster distributions compared to the classical k-means (round clusters) and Gaussian mixture models (ellipsoidal clusters). Given a dataset $X:=[x_{1}|\dots|x_{n}]\in\mathbb{R}^{p\times n}$, spectral clustering roughly consists in the eigenvalue decomposition of a Laplacian matrix $L\in\operatorname{Sym}_{n}$ associated to a pairwise similarity matrix, for instance $M_{ij}=\exp(\left\lVert x_{i}-x_{j}\right\rVert^{2}/2\sigma^{2})$. The eigenvectors are then used as new embeddings for the data points, on which standard clustering algorithms like k-means can be performed. This method is closely related to the celebrated Laplacian eigenmaps (Belkin and Niyogi, 2003) which are used for nonlinear dimension reduction. The good performances of such a simple method as spectral clustering are theoretically justified by the particular structure of the Laplacian matrix $L$ in an ideal case—block-diagonal with a multiple eigenvalue related to the number of clusters (Ng et al., 2001). The recent sparse spectral clustering (SSC) method (Lu et al., 2016) builds on such an ideal case and encourages the block-diagonality by looking for a sparse and low-rank representation:
$\underset{U\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\langle UU^{\top},L\rangle_{F}+\beta\left\lVert UU^{\top}\right\rVert_{1},$ (20)
@@ 4.4.1 Application of the flag trick to spectral clustering
The Grassmannian spectral clustering method (Wang et al., 2017) directly optimizes (20) on Grassmann manifolds:
$\underset{\mathcal{S}\in\operatorname{Gr}(p,q)}{\operatorname{argmin}}\,\langle\Pi_{\mathcal{S}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}}\right\rVert_{1}.$ (21)
The authors use a Riemannian trust region method (Absil et al., 2009) for optimization and show the interest of directly optimizing over the Grassmann manifold instead of convexifying the optimization space (Lu et al., 2016). The following proposition applies the flag trick to SSC.
@@
The flag trick applied to SSC (21) reads
$\underset{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\operatorname{argmin}}\,\langle\Pi_{\mathcal{S}_{1:d}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}_{1:d}}\right\rVert_{1}.$ (22)
and is equivalent to the following optimization problem:
$\underset{U_{1:d}\in\operatorname{St}(p,q)}{\operatorname{argmin}}\,\sum_{k=1}^{d}(d-(k-1))\operatorname{tr}\left({U_{k}}^{\top}LU_{k}\right)+\beta\left\lVert\sum_{k=1}^{d}(d-(k-1))U_{k}{U_{k}}^{\top}\right\rVert_{1}.$ (23)
Proof  The proof is given in Appendix (Appendix C).
We can see that the case $\beta=0$ corresponds to classical spectral clustering. Indeed, with a similar reasoning as in the proof of Theorem 4, we can easily show that the solution to problem (22) is explicit and corresponds to the flag of nested eigenspaces of $L$ (for increasing eigenvalues). Therefore, initializing the algorithm with the smallest $q$ eigenvectors of $L$ seems like a natural idea. Moreover, one may intuitively analyze the relative weighting of the mutually-orthogonal subspaces $\operatorname{Span}(U_{1})\perp\dots\perp\operatorname{Span}(U_{d})$ in (23) as favoring a model with $q_{1}$ clusters, and then adding successively $q_{k}-q_{k_{1}}$ clusters to improve the modeling of the Laplacian matrix, with a tradeoff between too much and too few clusters.

### v2

@@ 4.3.1 Application of the Flag Trick to Trace Ratio Problems
$\operatorname*{arg\,max}_{\mathcal{S}\in\operatorname{Gr}(p,q)}\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}}B\right)}.$ (17)
@@
The flag trick transforms the trace ratio problem (17) into:
$\operatorname*{arg\,max}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}\frac{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)}{\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)},$ (18)
which is equivalent to the following optimization problem:
$\operatorname*{arg\,max}_{U_{1:d}\in\operatorname{St}(p,q)}\frac{\sum_{k=1}^{d}(d-(k-1))\operatorname{tr}\left({U_{k}}^{\top}A{U_{k}}\right)}{\sum_{l=1}^{d}(d-(l-1))\operatorname{tr}\left({U_{l}}^{\top}B{U_{l}}\right)}.$ (19)
Proof The proof is given in Appendix B.
Equation (19) tells us several things. First, the subspaces $\penalty 10000\ {\operatorname{Span}(U_{1})\perp\dots\perp\operatorname{Span}(U_{d})}$ are weighted decreasingly, which means that they have less and less importance with respect to the TR objective. Second, we can see that the nested trace ratio problem (18) somewhat maximizes the numerator $\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}A\right)$ while minimizing the denominator $\operatorname{tr}\left(\Pi_{\mathcal{S}_{1:d}}B\right)$. Both subproblems have an explicit solution corresponding to our nested PCA Theorem 8. Hence, one can naturally initialize the steepest descent algorithm with the $q$ highest eigenvalues of $A$ or the $q$ lowest eigenvalues of $B$ depending on the application. For instance, for LDA, initializing Algorithm 1 with the highest eigenvalues of $A$ would spread the classes far apart, while initializing it with the lowest eigenvalues of $B$ would concentrate the classes, which seems less desirable since we do not want the classes to concentrate at the same point.
@@ 4.3.2 Nestedness Experiments for Trace Ratio Problems
We run some nestedness and classification experiments for the specific Trace Ratio problem of LDA. Many other applications (marginal Fisher analysis (yan_graph_2007), local discriminant embedding (chen_local_2005), etc.) could be similarly investigated.
First, we consider a synthetic dataset with five clusters. The ambient dimension is $p=3$ and the intrinsic dimensions that we try are $q_{1:2}=(1,2)$. We adopt a preprocessing strategy similar to ngo_trace_2012: we first center the data, then run a PCA to reduce the dimension to nCn-C (if nC<pn-C<p), then construct the LDA scatter matrices $A$ and $B$, then add a diagonal covariance regularization of $10^{-5}$ times their trace and finally normalize them to have unit trace. We run Algorithm 1 on Grassmann manifolds to solve the TR maximization problem (17), successively for $q_{1}=1$ and $q_{2}=2$. Then we plot the projections of the data points onto the optimal subspaces. We compare them to the nested projections onto the optimal flag output by running Algorithm 1 on $\operatorname{Fl}(3,(1,2))$ to solve (18). The results are shown in Figure 5. We can see that the Grassmann representations are non-nested while their flag counterparts perfectly capture the filtration of subspaces that best and best approximates the distribution while discriminating the classes. Even if the colors make us realize that the issue in this experiment for LDA is not much about the non-nestedness but rather about the rotation of the principal axes within the 2D subspace, we still have an important issue of consistency.
@@
Second, we consider the (full) handwritten digits dataset (alpaydin_optical_1998). It contains $8\times 8$ pixels images of handwritten digits, from 0 to $9$, almost uniformly class-balanced. One has $n=1797$, $p=64$ and $C=10$. We run a steepest descent algorithm to solve the trace ratio problem (18). We choose the full signature $q_{1:d}=(1,2,\dots,63)$ with $d=63$ and compare the output flag to the individual subspaces output by running optimization on $\operatorname{Gr}(p,q_{k})$ for $q_{k}\in q_{1:d}$. We plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ and the explained variance ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ as a function of the $k$. The results are illustrated in Figure 6. We see that the subspace angles are always positive and even very large sometimes with the LDA. Worst, the explained variance is not monotonic. This implies that we sometimes lose some information when increasing the dimension, which is extremely paradoxical.
Figure 6: Illustration of the nestedness issue in linear discriminant analysis (trace ratio problem) on the digits dataset. For $q_{k}\in(1,2,\dots,63)$, we solve the Grassmannian optimization problem (17) on $\operatorname{Gr}(64,q_{k})$ and plot the subspace angles $\Theta(\mathcal{S}_{k}^{*},\mathcal{S}_{k+1}^{*})$ (left) and explained variances ${\operatorname{tr}(\Pi_{\mathcal{S}_{k}^{*}}XX^{\top})}/{\operatorname{tr}(XX^{\top})}$ (right) as a function of $k$. We compare those quantities to the ones obtained by solving the flag optimization problem (18). We can see that the Grassmann-based method is highly non-nested and even yields an extremely paradoxical non-increasing explained variance (cf. red circle on the right).
@@
Third, we perform some classification experiments on the optimal subspaces. For each dataset, we run the optimization problems on $\operatorname{Fl}(p,q_{1:d})$, then project the data onto the different subspaces in $\mathcal{S}_{1:d}^{*}$ and run a nearest neighbors classifier with $5$ neighbors. The predictions are then ensembled (cf. Algorithm 2) by weighted averaging, either with uniform weights or with weights minimizing the average cross-entropy:
$w_{1}^{*},\dots,w_{d}^{*}=\operatorname*{arg\,min}_{\begin{subarray}{c}w_{k}\geq 0\\ \sum_{k=1}^{d}w_{k}=1\end{subarray}}-\frac{1}{nC}\sum_{i=1}^{n}\sum_{c=1}^{C}y_{ic}\ln\left(\sum_{k=1}^{d}w_{k}y_{kic}^{*}\right),$ (20)
@@
where $y_{kic}^{*}\in[0,1]$ is the predicted probability that $x_{i}\in\mathbb{R}^{p}$ belongs to class $c\in\{1,\dots,C\}$, by the classifier $g_{k}^{*}$ that is trained on $Z_{k}\coloneqq{U_{k}^{*}}^{\top}X\in\mathbb{R}^{q_{k}\times n}$. One can show that the latter is a convex problem, which we optimize using the cvxpy Python package (diamond2016cvxpy). We repeat the experiment 10 times in a stratified train-test fashion and report the average test cross-entropy in Table 1.
@@
Table 1: Results of the TR-based classification experiment. For each method (Gr: Grassmann optimization (17), Fl: flag optimization (18), Fl-U: flag optimization + uniform soft voting, Fl-W: flag optimization + optimal soft voting (20)), we give the average test cross-entropy between the predictions and the true labels.
@@
digits $1797$ $64$ $(1,2,5,10)$ $5.1$ $4.6$ $3.1$ $2.9$ $(0.03,0.05,0.30,0.62)$
wine $178$ $13$ $(1,2,5)$ $0.71$ $0.69$ $0.43$ $0.29$ $(0,0.68,0.32)$
breast $569$ $30$ $(1,2,5)$ $0.534$ $0.537$ $0.485$ $0.475$ $(0.26,0.24,0.50)$
iris $150$ 44 $(1,2,3)$ $0.275$ $0.271$ $0.281$ $0.265$ $(0.27,0.12,0.62)$
The wine example is particularly interesting. It first tells us that the optimal 5D subspace obtained by Grassmann optimization discriminates less between the classes than the 5D subspace from the optimal flag. This may show that the flag takes into account some lower dimensional variability that enables to better discriminate the classes. Second, the uniform averaging of the predictors at different dimensions improves the classification. Third, the optimal weights improve even more the classification and tell that the best discrimination occurs by taking a soft blend of classifiers at dimensions $2$ and $5$. Similar kinds of analyses can be made for the other examples.
@@ 4.3.3 Discussion on TR Optimization and Kernelization
To substantiate the generality of the flag trick, we propose two extensions of the previously proposed multilevel TR method: a Newton optimization algorithm and a kernelization. The experimentation of these extensions is left to future work since the goal is rather to show how the flag trick principle can be easily applied in different contexts.
@@ A Newton algorithm
In all the experiments of this paper, we use a steepest descent method on flag manifolds (Algorithm 1) to solve the flag problems. However, for the specific problem of TR (18), we believe that more adapted algorithms should be derived to take into account the specific form of the objective function, which is classically solved via a Newton-Lanczos method (ngo_trace_2012). To that extent, we develop in the appendix (B) an extension of the baseline Newton-Lanczos algorithm for the flag-tricked problem (18). In short, the latter can be reformulated as a penalized optimization problem of the form $\operatorname{argmax}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(p,q_{1:d})}{\sum_{k=1}^{d}\operatorname{tr}\left(\Pi_{\mathcal{S}_{k}}(A-\rho B)\right)}$, where $\rho$ is updated iteratively according to a Newton scheme. Once again, our central Theorem 8 enables to get an explicit solution to the penalized optimization problem.
@@
The classical trace ratio problems look for linear embeddings of the data. However, in most cases, the data follow a nonlinear distribution, which may cause linear dimension reduction methods to fail. The kernel trick (hofmann_kernel_2008) is a well-known method to embed nonlinear data into a linear space and fit linear machine learning methods. As a consequence, we propose in appendix (B) a kernelization of the trace ratio problem (18) in the same fashion as the one of the seminal graph embedding work (yan_graph_2007). This is expected to yield much better embedding and classification results.
@@ 4.4 The Flag Trick for Spectral Clustering Problems
As shown in introduction and particularly in Figure 1, spectral clustering methods based on Grassmannian optimization might suffer from the nestedness issue, which notably yields inconsistent data representations across different dimensions. This can be seen on the right part of the figure, where the clusters are almost separable in 1D but surprisingly intertwine in 2D. In this subsection, we address the nestedness issue by applying the flag trick (Definition 10) to Grassmann-based sparse spectral clustering (wang_grassmannian_2017).
Spectral clustering (ng_spectral_2001) is a baseline clustering technique. It can be applied to general cluster distributions compared to the classical k-means (round clusters) and Gaussian mixture models (ellipsoidal clusters). Given a dataset $X\coloneqq\begin{bmatrix}x_{1}&\cdots&x_{n}\end{bmatrix}\in\mathbb{R}^{p\times n}$, spectral clustering roughly consists in the eigenvalue decomposition of a Laplacian matrix $L\in\operatorname{Sym}(n)$ associated to a pairwise similarity matrix, for instance $M_{ij}=\exp(\left\lVert x_{i}-x_{j}\right\rVert^{2}/2\sigma^{2})$. The eigenvectors are then used as new embeddings for the data points, on which standard clustering algorithms like k-means can be performed. This method is closely related to the celebrated Laplacian eigenmaps (belkin_laplacian_2003) which are used for nonlinear dimension reduction. The good performances of such a simple method as spectral clustering are theoretically justified by the particular structure of the Laplacian matrix $L$ in an ideal case—block-diagonal with a multiple eigenvalue related to the number of clusters (ng_spectral_2001). The recent sparse spectral clustering (SSC) method (lu_convex_2016) builds on such an ideal case and encourages the block-diagonality by looking for a sparse and low-rank representation:
$\operatorname*{arg\,min}_{U\in\operatorname{St}(n,q)}\langle UU^{\top},L\rangle_{F}+\beta\left\lVert UU^{\top}\right\rVert_{1},$ (21)
@@ 4.4.1 Application of the Flag Trick to Spectral Clustering
The Grassmannian spectral clustering method (wang_grassmannian_2017) directly optimizes (21) on Grassmann manifolds:
$\operatorname*{arg\,min}_{\mathcal{S}\in\operatorname{Gr}(n,q)}\langle\Pi_{\mathcal{S}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}}\right\rVert_{1}.$ (22)
The authors use a Riemannian trust region method (absil_optimization_2008) for optimization and show the interest of directly optimizing over the Grassmann manifold instead of convexifying the optimization space (lu_convex_2016). The following proposition applies the flag trick to SSC.
@@
The flag trick transforms the sparse spectral clustering problem (22) into:
$\operatorname*{arg\,min}_{\mathcal{S}_{1:d}\in\operatorname{Fl}(n,q_{1:d})}\langle\Pi_{\mathcal{S}_{1:d}},L\rangle_{F}+\beta\left\lVert\Pi_{\mathcal{S}_{1:d}}\right\rVert_{1},$ (23)
which is equivalent to the following optimization problem:
$\operatorname*{arg\,min}_{U_{1:d}\in\operatorname{St}(n,q)}\sum_{k=1}^{d}(d-(k-1))\operatorname{tr}\left({U_{k}}^{\top}LU_{k}\right)+\beta\left\lVert\sum_{k=1}^{d}(d-(k-1))U_{k}{U_{k}}^{\top}\right\rVert_{1}.$ (24)
Proof The proof is given in Appendix C.
We can see that the case $\beta=0$ corresponds to classical spectral clustering. Indeed, with a similar reasoning as in the proof of Theorem 8, we can easily show that the solution to problem (23) is explicit and corresponds to the flag of nested eigenspaces of $L$ (for increasing eigenvalues). Therefore, initializing the algorithm with the smallest $q$ eigenvectors of $L$ seems like a natural idea. Moreover, one may intuitively analyze the relative weighting of the mutually-orthogonal subspaces $\operatorname{Span}(U_{1})\perp\dots\perp\operatorname{Span}(U_{d})$ in (24) as favoring a model with $q_{1}$ clusters, and then adding successively $q_{k}-q_{k-1}$ clusters to improve the modeling of the Laplacian matrix, with a tradeoff between too much and too few clusters.